# Norming Markushevich bases: recent results and open problems

### v1

@@ 6. The heredity problem
In the previous section we saw Alexandrov–Plichko’s theorem (Theorem 5.5) that the Banach space $\mathcal{C}([0,\omega_{1}])$ does not admit a norming M-basis and its generalisation from [85] that it does not embed in a Banach space with norming M-basis. While the latter result is obviously formally stronger, it is apparently not known whether admitting a norming M-basis and embedding in a Banach space with norming M-bases are distinct conditions in general. In other words, to the best of the authors’ knowledge, the following problem is still open.
@@
To begin with, it follows from Theorem 4.3 that the existence of a shrinking M-basis passes to subspaces (because both the WLD and the Asplund classes are hereditary). On the other hand, a result of Plichko [82, Theorem 2] (see, e.g., [31, Theorem 4.63]) claims that the dual of every separable Banach space embeds in a space with an M-basis; in particular, $\ell_{\infty}$ is a subspace of a Banach space with M-basis. Therefore, the existence of M-bases does not pass to subspaces.
@@
In contrast to Plichko’s result, $\ell_{\infty}$ does not embed in a Banach space with a norming M-basis. Indeed, as we saw in Section 4, Banach spaces with norming M-basis admit an LUR norm, while $\ell_{\infty}$ doesn’t have such a norm [23, Theorem II.7.10]
@@
As stated, the result appeared for the first time in [100, Proposition 4.6] and [36, Lemma 2.2]. A weaker result, under the assumption that $\mathcal{X}$ is WCG, was proved in [57, Proposition 6]; Vanderwerff [99, Proposition 3.1] proved the stronger claim that the $\lambda$-norming M-basis of $\mathcal{Z}$ can be extended to a $\lambda$-norming M-basis of $\mathcal{X}$. We sketch the proof, following [36].
@@
Conditions (6.1) and (6.2) then imply the existence of an M-basis $\{\tilde{u}_{\beta};\tilde{\psi}_{\beta}\}_{\beta\in\Lambda}$ in $\mathcal{Z}$ with the property that ${\rm span}\{\tilde{u}_{\beta}\}_{\beta\in\Lambda}={\rm span}\{u_{\beta}\}_{\beta\in\Lambda}$ and ${\rm span}\{\tilde{\psi}_{\beta}\}_{\beta\in\Lambda}={\rm span}\{\psi_{\beta}\}_{\beta\in\Lambda}$; in particular, ${\rm span}\{\tilde{\psi}_{\beta}\}_{\beta\in\Lambda}$ is $1$-norming, as desired. The proof of the existence of such M-basis uses the well-known method to construct projectional resolutions of the identity by transfinite induction. However, its explanation would require too long of a digression, hence we omit the argument and we refer to [100, Theorem 2.3] for the proof. ∎
@@
In connection to this problem, let us recall that is not even known if subspaces of $\ell_{1}(\Gamma)$ are Plichko, which was asked by Kalenda in [61, Question 4.44]. Notice that hyperplanes, as well as closed sublattices, of $\ell_{1}(\Gamma)$ are $1$-Plichko, [61, Example 4.39, Corollary 6.11].
We now pass to the second part of the section where we briefly discuss the analogue to Problem 6.1 for countably norming M-bases, or equivalently the heredity problem for Plichko spaces, that was asked for instance in [61, Question 4.45].
@@
Differently from the problem for norming M-bases, this problem has been solved and a negative answer follows from Kubiś’ results, [66, 67]. The solution is based upon two ingredients. The first one, of topological nature, is the result from [66, Theorem 5.8], containing the construction of a 00-dimensional linearly ordered Valdivia compact space $\mathcal{V}_{\omega_{1}}$ such that every linearly ordered Valdivia compactum is order preserving image of $\mathcal{V}_{\omega_{1}}$ (the notation $\mathcal{V}_{\omega_{1}}$ actually comes from [67]); here it is perhaps worth noting that linearly ordered Valdivia compacta have weight at most $\omega_{1}$, [66, Proposition 5.5]. Then $\mathcal{C}(\mathcal{V}_{\omega_{1}})$ is $1$-Plichko, by [61, Theorem 5.2]. Now, consider the ‘connectification’ $\mathcal{K}_{\omega_{1}}$ of $\mathcal{V}_{\omega_{1}}$ (see [67, Section 5] for the definition), which is an increasing continuous image of $\mathcal{V}_{\omega_{1}}$. Therefore, it follows that $\mathcal{C}(\mathcal{K}_{\omega_{1}})$ is a subspace of $\mathcal{C}(\mathcal{V}_{\omega_{1}})$. Finally, the last ingredient is the fact that $\mathcal{C}(\mathcal{K}_{\omega_{1}})$ is not Plichko, [67, Theorem 5.1]. Therefore, Plichko spaces are not closed under taking subspaces.
@@
At this point, it is of course natural to ask whether $\mathcal{C}(\mathcal{V}_{\omega_{1}})$ could admit a norming M-basis, since this would solve Problem 6.1. However, by the very property of $\mathcal{V}_{\omega_{1}}$, $[0,\omega_{1}]$ is a continuous image of $\mathcal{V}_{\omega_{1}}$, therefore $\mathcal{C}([0,\omega_{1}])$ is a subspace of $\mathcal{C}(\mathcal{V}_{\omega_{1}})$. Hence the latter doesn’t admit a norming M-basis, due to Theorem 5.8.
However, we can actually say more. In fact, another property of $\mathcal{V}_{\omega_{1}}$ is that it is ‘saturated’ by copies of $[0,\omega_{1}]$, [66, Theorem 5.8]. Moreover, it is easy to see that every closed subset of a linearly ordered compactum admits a regular extension operator, [67, Lemma 4.2]. Therefore, $\mathcal{C}([0,\omega_{1}])$ is even $1$-complemented in $\mathcal{C}(\mathcal{V}_{\omega_{1}})$.
@@
A positive answer to this is known under the additional assumption that $\mathcal{X}/\mathcal{Z}$ is separable, [61, Theorem 4.40]. Moreover, Kubiś proved that every $1$-complemented subspace of a $1$-Plichko Banach space of density $\omega_{1}$ is $1$-Plichko as well, [67, Theorem 6.3].
@@
Notice that $\mathcal{C}([0,\eta])$ is not $1$-Plichko when $\eta\geqslant\omega_{2}$, because then $[0,\eta]$ is not Valdivia by [61, Example 1.10] and then the claim follows from [61, Theorem 5.3]; see also [65, Theorem 5.10].
One might wonder whether, similarly as Section 5, one could improve Kalenda’s result [62] by proving that $\mathcal{C}([0,\omega_{2}])$ does not embed in a Plichko space. At the moment, Kalenda’s proof doesn’t seem to give this stronger claim, therefore the following problem is also open.
@@ 7. Semi-Eberlein compacta
As we saw in Section 3, the map $\mathsf{ev}$ from (1.1) allows to characterise several classes of non-separable Banach spaces in terms of topological properties of the dual unit ball with the $w^{*}$-topology. In this section we focus on the class of compact spaces whose definition is obtained by describing the image of $\mathsf{ev}$ when the M-basis is $1$-norming. Such compacta were introduced in [69], where they are called semi-Eberlein compacta, and some results concerning them were recently obtained in [14, 16, 45]; still, their structure is rather poorly understood as of today.
@@
By the very definition, every Eberlein compactum is semi-Eberlein and every semi-Eberlein compactum is Valdivia. Moreover, every cube $[-1,1]^{\Gamma}$ is by definition semi-Eberlein and the ordinal interval $[0,\omega_{1}]$ is not semi-Eberlein, as we will discuss; thus these three classes are distinct. The table below illustrates the fact that semi-Eberlein compacta are obtained from Eberlein ones via the same generalisation that leads from Corson to Valdivia compacta.
@@
Let us start once more by considering $[0,\omega_{1}]$. In order to explain why this compact space is not semi-Eberlein, we need to recall some topological notions. Recall that a point $p$ in a topological space $\mathcal{T}$ is a P-point if $p$ is not isolated and for every countable family $(\mathcal{U}_{k})_{k=1}^{\infty}$ of neighbourhoods of $p$, $\cap_{k=1}^{\infty}\mathcal{U}_{k}$ is a neighbourhood of $p$. A point $p\in\mathcal{T}$ is a weak P-point if $p$ is not isolated and it is limit point of no countable set in $\mathcal{T}\setminus\{p\}$. An archetypal example of P-point is the point $\omega_{1}$ in $[0,\omega_{1}]$.
Kubiś and Leiderman proved in [69, Theorem 4.2] that semi-Eberlein compacta do not admit P-points, hence it follows that $[0,\omega_{1}]$ fails to be semi-Eberlein. Incidentally, this result can be combined with a forcing argument to give an example of a Corson compactum that is not semi-Eberlein, [69, Example 5.5]. It was then natural to ask whether semi-Eberlein compacta could admit weak P-points, [69, Question 6.1], which was answered in [45, Section 4.3], once more by consideration of $\mathcal{K}_{\varrho}$.
@@
Let us now turn to the question to study dual balls that are semi-Eberlein. For instance, the dual ball of $\ell_{1}(\Gamma)$ is semi-Eberlein, as we already noted. Moreover, when $\mathcal{X}$ is WCG (or, more generally, subspace of WCG), $(B_{\mathcal{X}^{*}},w^{*})$ is semi-Eberlein as well. In particular, by taking $\mathcal{X}$ to be the Banach space from Theorem 4.5, we see that $(B_{\mathcal{X}^{*}},w^{*})$ is semi-Eberlein and yet $\mathcal{X}$ doesn’t admit any norming M-basis. Concerning the opposite implication, we have the following remark.
@@
Notice that in Remark 7.3 we actually proved that $(B_{\mathcal{X}^{*}},w^{*})$ embeds in $[-1,1]^{\Gamma}$ in such a way that $B_{\mathcal{X}^{*}}\cap c_{00}(\Gamma)$ is dense in $(B_{\mathcal{X}^{*}},w^{*})$. Therefore it is natural to ask whether this behaviour of density of finitely supported vectors is common to all semi-Eberlein spaces; this problem was already asked in [69, Problem 6.4].
@@
Let $\mathcal{K}$ be a semi-Eberlein compactum. Is there a homeomorphic embedding $\varphi\colon\mathcal{K}\to[-1,1]^{\Gamma}$ such that $\varphi[\mathcal{K}]\cap c_{00}(\Gamma)$ is dense in $\varphi[\mathcal{K}]$?
@@
Notice that a positive answer to the problem has been recently given for $\mathcal{K}$ metrisable, [74, Proposition 6.5]. Concerning (1), recall that an Eberlein compactum is scattered if and only if it is strong Eberlein, namely it is homeomorphic to a compact subset of $c_{00}(\Gamma)\cap\{0,1\}^{\Gamma}$, [4]. This does not extend to semi-Eberlein compacta, since $\{0,1\}^{\Gamma}$ is not scattered, but $c_{00}(\Gamma)\cap\{0,1\}^{\Gamma}$ is clearly dense in $\{0,1\}^{\Gamma}$.
To conclude, let us point out that, as a rule, techniques used to prove results for Valdivia or Eberlein compacta do not admit obvious modifications for semi-Eberlein compacta. Perhaps the unique exception is a Rosenthal-type characterisation of semi-Eberlein compacta, [69, Proposition 3.1], analogous to [83, Theorem 3.1] (see also [61, Proposition 1.9]). For instance, it is not known if semi-Eberlein compacta admit a similar decomposition as the one obtained by Farmaki [33] for Eberlein compacta, see [14, Question 47]. Very little is also known concerning continuous images of semi-Eberlein compacta; the most interesting result here is perhaps that the class of Corson semi-Eberlein compacta is stable under continuous images, [14, Theorem C] (notice that the said class is strictly larger than that of Eberlein compacta, [69, Section 3], [91]). Some other explicit questions can be found in [69, Section 6] (note that Question 1 and the second part of Question 6 there have been answered in [45] and [14] respectively).
Here we limit ourselves to asking the following problem, analogue of a well-known open question due to Avilés and Kalenda [9, Problem 16] for Valdivia compacta.
@@
Acknowledgements. The authors wish to express their gratitude to Antonio Avilés, Marián Fabian, Gilles Godefroy, Chris Lambie-Hanson, Grzegorz Plebanek, and Jacopo Somaglia for several helpful remarks on various parts of the paper.

### v2

@@ 6. The heredity problem
In the previous section a central role was played by Alexandrov–Plichko’s theorem (Theorem 5.4) that the Banach space $\mathcal{C}([0,\omega_{1}])$ does not admit a norming M-basis and its generalisation from [96] that it does not embed in a Banach space with norming M-basis. While the latter result is obviously formally stronger, it is apparently not known whether admitting a norming M-basis and embedding in a Banach space with norming M-bases are distinct conditions in general. In other words, to the best of the authors’ knowledge, the following problem is still open.
@@
To begin with, it follows from Theorem 4.3 that the existence of a shrinking M-basis passes to subspaces (because both the WLD and the Asplund classes are hereditary). On the other hand, a result of Plichko [93, Theorem 2] (see, e.g., [38, Theorem 4.63]) claims that the dual of every separable Banach space embeds in a space with an M-basis; in particular, $\ell_{\infty}$ is a subspace of a Banach space with M-basis. Therefore, the existence of M-bases does not pass to subspaces.
@@
In contrast to Plichko’s result, $\ell_{\infty}$ does not embed in a Banach space with a norming M-basis. Indeed, as we saw in Section 4, Banach spaces with norming M-basis admit an LUR norm, while $\ell_{\infty}$ does not have such a norm [28, Theorem II.7.10]
@@
As stated, the result appeared for the first time in [111, Proposition 4.6] and [43, Lemma 2.2]. A weaker result, under the assumption that $\mathcal{X}$ is WCG, was proved in [64, Proposition 6]; Vanderwerff [110, Proposition 3.1] proved the stronger claim that the $\lambda$-norming M-basis of $\mathcal{Z}$ can be extended to a $\lambda$-norming M-basis of $\mathcal{X}$. We sketch the proof, following [43].
@@
Conditions (6.1) and (6.2) then imply the existence of an M-basis $\{x_{\beta};f_{\beta}\}_{\beta\in\Lambda}$ in $\mathcal{Z}$ with the property that ${\rm span}\{x_{\beta}\}_{\beta\in\Lambda}={\rm span}\{u_{\beta}\}_{\beta\in\Lambda}$ and ${\rm span}\{f_{\beta}\}_{\beta\in\Lambda}={\rm span}\{g_{\alpha}\}_{\alpha\in\Gamma}$; in particular, ${\rm span}\{f_{\beta}\}_{\beta\in\Lambda}$ is $1$-norming, as desired. The proof of the existence of such an M-basis uses the well-known method to construct projectional resolutions of the identity by transfinite induction. However, its explanation would require too long of a digression, hence we omit the argument and we refer to [111, Theorem 2.3] for the proof. ∎
@@
In connection to this problem, let us recall that is not even known if subspaces of $\ell_{1}(\Gamma)$ are Plichko, which was asked by Kalenda in [70, Question 4.44]. Notice that hyperplanes, as well as closed sublattices, of $\ell_{1}(\Gamma)$ are $1$-Plichko, [70, Example 4.39, Corollary 6.11].
We now pass to the second part of the section where we briefly discuss the analogue to Problem 6.1 for countably norming M-bases, or equivalently the heredity problem for Plichko spaces, that was asked for instance in [70, Question 4.45].
@@
Differently from the problem for norming M-bases, this problem has been solved and a negative answer follows from Kubiś’ results, [75, 76]. The solution is based upon two ingredients. The first one, of topological nature, is the result from [75, Theorem 5.8], containing the construction of a 00-dimensional linearly ordered Valdivia compact space $\mathcal{V}_{\omega_{1}}$ such that every linearly ordered Valdivia compactum is order preserving image of $\mathcal{V}_{\omega_{1}}$ (the notation $\mathcal{V}_{\omega_{1}}$ actually comes from [76]); here it is perhaps worth noting that linearly ordered Valdivia compacta have weight at most $\omega_{1}$, [75, Proposition 5.5]. Then $\mathcal{C}(\mathcal{V}_{\omega_{1}})$ is $1$-Plichko, by [70, Theorem 5.2]. Now, consider the ‘connectification’ $\mathcal{K}_{\omega_{1}}$ of $\mathcal{V}_{\omega_{1}}$ (see [76, Section 5] for the definition), which is a (two-to-one) order preserving continuous image of $\mathcal{V}_{\omega_{1}}$. Therefore, it follows that $\mathcal{C}(\mathcal{K}_{\omega_{1}})$ is a subspace of $\mathcal{C}(\mathcal{V}_{\omega_{1}})$. Finally, the last ingredient is the fact that $\mathcal{C}(\mathcal{K}_{\omega_{1}})$ is not Plichko, [76, Theorem 5.1]. Therefore, Plichko spaces are not closed under taking subspaces.
@@
At this point, it is of course natural to ask whether $\mathcal{C}(\mathcal{V}_{\omega_{1}})$ could admit a norming M-basis, since this would solve Problem 6.1. However, by the very property of $\mathcal{V}_{\omega_{1}}$, $[0,\omega_{1}]$ is a continuous image of $\mathcal{V}_{\omega_{1}}$, therefore $\mathcal{C}([0,\omega_{1}])$ is a subspace of $\mathcal{C}(\mathcal{V}_{\omega_{1}})$. Hence the latter does not admit a norming M-basis, due to Theorem 5.8.
However, we can actually say more. In fact, another property of $\mathcal{V}_{\omega_{1}}$ is that it is ‘saturated’ by copies of $[0,\omega_{1}]$, [75, Theorem 5.8]. Moreover, it is easy to see that every closed subset of a linearly ordered compactum admits a regular extension operator, [76, Lemma 4.2]. Therefore, $\mathcal{C}([0,\omega_{1}])$ is even $1$-complemented in $\mathcal{C}(\mathcal{V}_{\omega_{1}})$.
@@
A positive answer to this is known under the additional assumption that $\mathcal{X}/\mathcal{Z}$ is separable, [70, Theorem 4.40]. Moreover, Kubiś proved that every $1$-complemented subspace of a $1$-Plichko Banach space of density $\omega_{1}$ is $1$-Plichko as well, [76, Theorem 6.3].
@@
Notice that $\mathcal{C}([0,\eta])$ is not $1$-Plichko when $\eta\geqslant\omega_{2}$, because $[0,\eta]$ is not Valdivia for $\eta\geqslant\omega_{2}$ by [70, Example 1.10] and then the claim follows from [70, Theorem 5.3]; see also [74, Theorem 5.10].
One might wonder whether, similarly as Section 5, one could improve Kalenda’s result [71] by proving that $\mathcal{C}([0,\omega_{2}])$ does not embed in a Plichko space. At the moment, Kalenda’s proof does not seem to give this stronger claim, therefore the following problem is also open.
@@ 7. Semi-Eberlein compacta
As we saw in Section 3, the map $\mathsf{ev}$ from (1.1) allows one to characterise several classes of non-separable Banach spaces in terms of topological properties of the dual unit ball with the $w^{*}$-topology. In this section we focus on the class of compact spaces whose definition is obtained by describing the image of $\mathsf{ev}$ when the M-basis is $1$-norming. Such compacta were introduced in [78], where they are called semi-Eberlein compacta, and some results concerning them were recently obtained in [19, 21, 52]; still, their structure is rather poorly understood as of today.
@@
By the very definition, every Eberlein compactum is semi-Eberlein and every semi-Eberlein compactum is Valdivia. Moreover, every cube $[-1,1]^{\Gamma}$ is by definition semi-Eberlein and the ordinal interval $[0,\omega_{1}]$ is not semi-Eberlein, as we will discuss; thus these three classes are distinct. The table below illustrates the fact that semi-Eberlein compacta are obtained from Eberlein compacta via the same generalisation that leads from Corson to Valdivia compacta.
@@
Let us start once more by considering $[0,\omega_{1}]$. In order to explain why this compact space is not semi-Eberlein, we need to recall some topological notions. Recall that a point $p$ in a topological space $\mathcal{T}$ is a P-point if $p$ is not isolated and for every countable family $(\mathcal{U}_{k})_{k=1}^{\infty}$ of neighbourhoods of $p$, $\cap_{k=1}^{\infty}\mathcal{U}_{k}$ is a neighbourhood of $p$. A point $p\in\mathcal{T}$ is a weak P-point if $p$ is not isolated and it is not a limit point of any countable set in $\mathcal{T}\setminus\{p\}$. An archetypal example of P-point is the point $\omega_{1}$ in $[0,\omega_{1}]$.
Kubiś and Leiderman proved in [78, Theorem 4.2] that semi-Eberlein compacta do not admit P-points, whence it follows that $[0,\omega_{1}]$ fails to be semi-Eberlein. Incidentally, this result can be combined with a forcing argument to give an example of a Corson compactum that is not semi-Eberlein, [78, Example 5.5]. It was then natural to ask whether semi-Eberlein compacta could admit weak P-points, [78, Question 6.1], which was answered in [52, Section 4.3], once more by consideration of $\mathcal{K}_{\varrho}$.
@@
Let us now turn to the consideration of semi-Eberlein dual balls. For instance, the dual ball of $\ell_{1}(\Gamma)$ is semi-Eberlein, as we already noted. Moreover, when $\mathcal{X}$ is WCG (or, more generally, subspace of WCG), $(B_{\mathcal{X}^{*}},w^{*})$ is semi-Eberlein as well. In particular, by taking $\mathcal{X}$ to be the Banach space from Theorem 4.5, we see that $(B_{\mathcal{X}^{*}},w^{*})$ is semi-Eberlein and yet $\mathcal{X}$ does not admit any norming M-basis. Concerning the opposite implication, we have the following remark.
@@
Notice that in Remark 7.3 we actually proved that $(B_{\mathcal{X}^{*}},w^{*})$ embeds in $[-1,1]^{\Gamma}$ in such a way that $B_{\mathcal{X}^{*}}\cap c_{00}(\Gamma)$ is dense in $(B_{\mathcal{X}^{*}},w^{*})$. Therefore it is natural to ask whether this behaviour of density of finitely supported vectors is common to all semi-Eberlein spaces; this problem was already asked in [78, Problem 6.4].
@@
Let $\mathcal{K}$ be a semi-Eberlein compactum. Is there a homeomorphic embedding $\varphi\colon\mathcal{K}\to[-1,1]^{\Gamma}$ such that $\varphi(\mathcal{K})\cap c_{00}(\Gamma)$ is dense in $\varphi(\mathcal{K})$?
@@
Notice that a positive answer to the problem has been recently given for $\mathcal{K}$ metrisable, [84, Proposition 6.5]. Concerning (1), recall that an Eberlein compactum is scattered if and only if it is strong Eberlein, namely it is homeomorphic to a compact subset of $c_{00}(\Gamma)\cap\{0,1\}^{\Gamma}$, [4]. This does not extend to semi-Eberlein compacta, since $\{0,1\}^{\Gamma}$ is not scattered, but $c_{00}(\Gamma)\cap\{0,1\}^{\Gamma}$ is clearly dense in $\{0,1\}^{\Gamma}$.
To conclude, let us point out that, as a rule, techniques used to prove results for Valdivia or Eberlein compacta do not admit obvious modifications for semi-Eberlein compacta. Perhaps the unique exception is a Rosenthal-type characterisation of semi-Eberlein compacta, [78, Proposition 3.1], analogous to [94, Theorem 3.1] (see also [70, Proposition 1.9]). For instance, it is not known if semi-Eberlein compacta admit a similar decomposition as the one obtained by Farmaki [40] for Eberlein compacta, see [19, Question 47]. Very little is also known concerning continuous images of semi-Eberlein compacta; the most interesting result here is perhaps that the class of Corson semi-Eberlein compacta is stable under continuous images, [19, Theorem C] (notice that the said class is strictly larger than that of Eberlein compacta, [78, Section 3], [102]). Some other explicit questions can be found in [78, Section 6] (note that Question 1 and the second part of Question 6 there have been answered in [52] and [19] respectively).
Here we limit ourselves to asking the following problem, which is an analogue of a well-known open question due to Avilés and Kalenda [9, Problem 16] for Valdivia compacta.
@@
Acknowledgements. The authors wish to express their gratitude to Antonio Avilés, Marián Fabian, Gilles Godefroy, Chris Lambie-Hanson, Gilles Lancien, Grzegorz Plebanek, and Jacopo Somaglia for several helpful remarks on various parts of the paper.